# by Leveraging CLIP Latent Space

### 07 Feb 2024 v1

@@ 3.2 Image-text Interleaved Training
Our approach targets developing a versatile prior model capable of processing diverse inputs to estimate visual outputs. Drawing from earlier methodologies, a straightforward solution involves concatenating different inputs, like combining text (“a dog wearing sunglasses”) with respective concept-specific images. However, preliminary experiments indicated that this method does not effectively capture the intricate relationships between target text tokens (e.g., “dog”) and the corresponding concept images.
@@
To address this, we adopt the interleaved pre-training strategy used in Kosmos-G, but with a notable modification to enhance resource efficiency. We incorporate pretrained CLIP text and vision encoders for extracting modality-specific embeddings—separating text-only from subject-specific images. The key refinement in our process is the substitution of subject token embeddings with corresponding vision embeddings instead of introducing additional tokens to handle the image embeddings via resampler [2]. This alteration allows us to bypass the need to train the prior or MLLMs, significantly improving the model’s proficiency in handling interleaved data.
@@
For the generation of high-quality training datasets, we carefully selected 2 million high-quality images from the LAION dataset [41], each with a resolution of 1024x1024. Utilizing BLIP2, we generate captions for these images and employ SAM [19] for extracting noun or subject-specific segmentation masks. Given the CLIP model’s requirement for 224x224 resolution images, we avoid resizing the masks within their original resolutions. Instead, we opt for cropping the area of interest using DINO [24], followed by resizing the masked object while preserving its aspect ratio. This technique is crucial in retaining maximum visual information for each subject during the training phase. We provide more details about the filters used in the supplementary material.
In summary, the ECLIPSE  prior, trained with our image-text interleaved data and supplementary conditions, presents an efficient pathway for P-T2I. The culmination of this process yields the $\lambda$-ECLIPSE model, a refined version characterized by enhanced effectiveness in handling and interpreting multimodal data.
@@ 4 Experiments
In this section, we first introduce the experimental setup, evaluations, and baselines. Later, we delve into the qualitative and quantitative results.
@@
We initialize our model, $\lambda$-ECLIPSE, equipped with 34M parameters. We train our model on an image-text interleaved dataset of 2M instances, partitioned into 1.6M for training and 0.4M for validation. The model is specifically tuned for the Kandinsky v2.2 diffusion image decoder. Therefore, we use pre-trained OpenCLIP-ViT-G/14222https://huggingface.co/laion/CLIP-ViT-g-14-laion2B-s12B-b42K as the text and vision encoders, ensuring alignment with Kandinsky v2.2 image embeddings. Training is executed on 2 x A100 GPUs, leveraging a per-GPU batch size of 512 and a peak learning rate of 0.00005, across approximately 100,000 iterations, summing up to 74 GPU hours. During inference, the model employs 50 DDIM steps and 7.5 classifier-free guidance for the Kandinsky v2.2 diffusion image generator. Adhering to baseline methodologies, we perform the P-T2I following the baseline papers’ protocols. For $\lambda$-ECLIPSE, target subject pixel regions in reference images are segmented before embedding extraction via the CLIP(vision) encoder. We drop the canny edge map during inference unless specified explicitly.
@@
We primarily utilize Dreambench (encompassing 30 unique concepts with 25 prompts per concept) for qualitative and quantitative evaluations using DINO and CLIP-based metrics [38]. Due to their limitations, we extend our evaluations on the ConceptBed [30] benchmark (covering 80 diverse imagenet concepts and a total of 33K composite prompts), where we report performance on concept replication, concept, and composition alignment using the Concept Confidence Deviation ($CCD(\downarrow)$) metric.
4.1 Results & Analysis
Table 2: Quantitative comparisons of different methodologies on Dreambench. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents that we reproduce the open-source methodologies. Other results are borrowed from the Subject-Diffusion [27].
Method Base Model DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
Textual Inversion SDv1.5 0.569 0.780 0.255
DreamBooth SDv1.5 0.668 0.803 0.305
Custom Diffusion SDv1.5 0.643 0.790 0.305
Re Imagen Imagen 0.600 0.740 0.270
ELITE SDv1.4 0.621 0.771 0.293
Subject-Diffusion SDv1.5 0.711 0.787 0.293
BLIP-Diffusion* SDv1.5 0.603 0.793 0.291
IP-Adapter* SDv1.5 0.629 0.827 0.264
IP-Adapter* SDXL 0.613 0.810 0.292
Kosmos-G* SDv1.5 0.618 0.822 0.250
@@ 4.1 Results & Analysis
Figure 6: Multi-subject qualitative examples illustrating the comparison between $\lambda$-ECLIPSE and other SOTA baselines.
Table 3: Quantitative comparisons of different methodologies on ConceptBed. We present results on $CCD~{}(\downarrow)$ across three evaluation categories. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents that we reproduce the open-source methodologies. Other results are borrowed from the ConceptBed [30].
@@
Model Replication Alignment Alignment
@@
Figure 7: Qualitative examples for edge-guided P-T2I showcasing the comparison between BLIP-Diffusion vs. $\lambda$-ECLIPSE.
@@
The quantitative assessments are detailed in Table 2 and Table 3, focusing on the single-concept T2I task. For Dreambench, we generate and evaluate four images per prompt, reporting average performance on three metrics (DINO, CLIP-I, and CLIP-T). In the case of ConceptBed, we process each of the 33K prompts to generate a single image. The results, as depicted in these tables, highlight $\lambda$-ECLIPSE’s superior performance in composition tasks and its competitive edge over even fine-tuning-based methodologies, while maintaining robust concept alignment. $\lambda$-ECLIPSE exhibits a notable proficiency in concept replication, albeit with a marginal trade-off in concept alignment for enhanced composition fidelity. Comparatively, IP-Adapter, ELITE, and BLIP-Diffusion prioritize concept alignment, often at the expense of composition alignment. Noteworthy is $\lambda$-ECLIPSE’s efficiency, achieved with significantly fewer resources.
@@
In Figure 5, we present a range of single subject-specific images generated by various methodologies including BLIP-Diffusion, IP-Adapter, Kosmos-G, and $\lambda$-ECLIPSE. $\lambda$-ECLIPSE demonstrates exemplary proficiency in composition while ensuring concept alignment. In contrast, the baselines often overemphasize reference images (notably in rows 2 and 4) or exhibit concept dilution (as seen in rows 6 and 8), leading to higher concept alignment but compromised composition. Additionally, Figure 6 exhibits $\lambda$-ECLIPSE’s multi-concept generation prowess, in comparison to ZipLoRA (fine-tuning-based approach) and Kosmos-G (Multimodal LLM-based approach), underscoring its capability to rival compute-intensive methodologies.
Figure 2 previously demonstrated Kandinsky v2.2’s ability to partially reconstruct canny edge maps while preserving semantic integrity. Building on this, Figure 7 offers qualitative insights, comparing BLIP-Diffusion and $\lambda$-ECLIPSE in edge-guided subject-driven T2I scenarios. Here, BLIP-Diffusion adheres strictly to the imposed conditions, often at the cost of concept retention (rows 1, 3, and 4). Conversely, $\lambda$-ECLIPSE not only complies with these conditions but also adeptly balances the need for high concept and composition alignment.
@@
Alignment Alignment
Projection loss 0.394 0.008
w/ contrastive loss (lambda=0.5) 0.435 -0.043
w/ contrastive loss (lambda=0.2) 0.402 -0.026
@@
We extend our study to evaluate the individual contributions of different components in $\lambda$-ECLIPSE. Initially, the model’s performance with solely the projection loss (referenced in Eq.1) is assessed. Subsequent experiments involve training $\lambda$-ECLIPSE variants with varying hyperparameters for the contrastive loss, specifically $\lambda$ values of 0.2 and 0.5. A comparative analysis of these baselines is conducted against the fully equipped $\lambda$-ECLIPSE model, which incorporates $\mathcal{L}_{prior}$ (Eq.1) with $\lambda=0.2$ and utilizes canny edge maps during training. As Table 4 illustrates, relying solely on projection loss results in high concept alignment but compromises compositions. The contrastive loss variant with $\lambda=0.5$ enhances composition alignment at the expense of concept alignment, whereas $\lambda=0.2$ achieves a more balanced performance. Crucially, the integration of canny edge maps during training optimally balances both alignments and, specifically, improves the concept alignment. Here, the negative values indicate that the $CCD$ oracle model performs similarly to real images having the same set of compositions.
@@
A key attribute of the CLIP latent space is the ability to perform smooth interpolation between two sets of embeddings. We conducted experiments to demonstrate $\lambda$-ECLIPSE’s ability to learn and replicate this smooth latent space transition. We selected two distinct dog breeds ($<$dog1$>$, $<$dog2$>$) and two types of hats ($<$hat1$>$, $<$hat2$>$) as the concepts. $\lambda$-ECLIPSE was then used to estimate image embeddings for all four possible combinations, each corresponding to the input phrase “a $<$dog$>$ wearing a $<$hat$>$.” Figure 8 showcases a gradual and seamless transition in the synthesized images from the top left to the bottom right. Unlike current diffusion models, which often exhibit sensitivity to input variations necessitating numerous iterations for desired outcomes, $\lambda$-ECLIPSE inherits CLIP’s smooth latent space. This inheritance not only facilitates progressive changes in the conceptual domain but also extends the model’s utility by enabling personalized multi-subject interpolations, a feature that is not inherently available in previous P-T2I methods.
@@ 5 Conclusion
In this paper, we have introduced a novel diffusion-free methodology for personalized text-to-image (T2I) applications, utilizing the latent space of the pre-trained CLIP model with high efficiency. Our $\lambda$-ECLIPSE model, trained on an image-text interleaved dataset of superior quality, marks a significant advancement in the field. This model outperforms existing methods by its capability to execute single-concept, multi-concept, and edge-guided controlled P-T2I tasks using a singular model framework, while simultaneously minimizing resource utilization. Notably, $\lambda$-ECLIPSE sets a new benchmark in achieving state-of-the-art results in composition alignment, coupled with exceptional maintenance of subject fidelity. Furthermore, our research illuminates the potential of $\lambda$-ECLIPSE in exploring and leveraging the smooth latent space. This capability unlocks new avenues for interpolating between multiple concepts and their amalgamation, thereby generating entirely novel concepts. Our findings underscore a promising pathway where pre-trained diffusion image generators can be effectively controlled without necessitating extra supervision.

### 09 Apr 2024 v2

@@ 3.2 Image-text Interleaved Training
Figure 4: This figure illustrates a qualitative comparison of $\lambda$-ECLIPSE  with contemporary approaches for single-subject T2I generations, utilizing concepts and prompts from the Dreambench dataset. For each method, concept, and prompt, we generate four images and select the one that most accurately represents the queried concept and composition.
@@
Our approach targets developing a versatile prior model capable of processing diverse inputs to estimate target visual outputs. Drawing from earlier methodologies, a straightforward solution involves concatenating different inputs, like combining text (“a dog wearing sunglasses”) with respective concept-specific images. Preliminary experiments indicated that this method does not effectively capture the intricate relationships between target text tokens (e.g. “dog”) and the corresponding concept images, especially when multiple concepts are present.
@@
To address this, we adopt the interleaved pre-training strategy used in Kosmos-G, but with a notable modification to enhance resource efficiency. We incorporate pretrained frozen CLIP text and vision encoders for extracting modality-specific embeddings—separating text-only from subject-specific images. The key refinement in our process is the substitution of subject token-specific text embeddings with corresponding vision embeddings instead of introducing additional trainable tokens to handle the image embeddings via resampler alayrac2022flamingo . First, we extract reference concept visual features ($z_{x_{k}}\in\mathcal{R}^{1\mathrm{x}1280}$) from the CLIP vision encoder. Similarly, we also extract the text prompt features ($z_{y}\in\mathcal{R}^{77\mathrm{x}1280}$) from the last layer of the CLIP text encoder. Here, 1280 is the CLIP-specific feature dimension. At last, we replace the concept noun corresponding latent features from $z_{y}$ with $z_{x_{k}}$; resulting in image-text interleaved features while preserving the contextual information of the text features. This alteration allows us to bypass the need to train the big priors models (e.g. MLLMs), significantly improving the model’s proficiency in handling interleaved data.
@@
For the generation of high-quality training datasets, we carefully selected 2 million high-quality images from the LAION dataset schuhmann2022laion , each with a resolution of 1024x1024. Utilizing BLIP2, we generate captions for these images and employ SAM kirillov2023segment for extracting noun or subject-specific segmentation masks. Given the CLIP model’s requirement for 224x224 resolution images, we avoid resizing the masks within their original resolutions. Instead, we opt for cropping the area of interest using Grounding DINO liu2023grounding , followed by resizing the masked object while preserving its aspect ratio. This technique is crucial in retaining maximum visual information for each subject during the training phase. We provide more details about the filters used in the supplementary material.
3.3 Additional Finetuning
Due to the nature of UnCLIP models, even if $\lambda$-ECLIPSE is very accurate, the diffusion UNet model ($h_{\phi}$) may not be effective in generating very unique visual representations. However, such behavior is common across the fast P-T2I methods and they lack in terms of maintaining performance compared to the finetuning-based methods (as outlined in Table 2). Therefore, we extend the $\lambda$-ECLIPSE and perform concept-specific finetuning.
Compared to the traditional finetuning methodologies (such as DreamBooth), $\lambda$-ECLIPSE provides very unique advantages. As $\lambda$-ECLIPSE prior model ($f_{\theta}$) is pre-trained for personalization, there is no need for further finetuning the $f_{\theta}$ and we need to only finetune diffusion UNet model. Importantly, the fine-tuning of the $h_{\phi}$ does not depend on the text embeddings ($z_{y}$). Hence, this leads to stable fine-tuning of the $h_{\phi}$; unlike DreamBooth on stable diffusion that observes catastrophic forgetting (Section D). The new fine-tuning objective is:
$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],(z_{x})\end{subarray}}{\mathbb{E}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x})||_{2}^{2}\Big{]}.$ (2)
Here, $z_{x}$ is the visual feature of the reference concept image $x$. Notably, we do not need to use regularization from the DreamBooth as text alignment is already ensured during the pretraining stage of $\lambda$-ECLIPSE. Moreover, this finetuning can be performed across the set of given visual concepts altogether in a single model without degrading performance.
In summary, the prior model, trained with our image-text interleaved data and supplementary condition, presents an efficient pathway for resource-efficient multi-subject-driven image generations.
@@ 4 Experiments
In this section, we first introduce the experimental and evaluation setups. Later, we delve into the qualitative and quantitative results.
@@
We initialize our model, $\lambda$-ECLIPSE, equipped with 34M parameters. We train our model on an image-text interleaved dataset of 2M instances, partitioned into 1.6M for training and 0.4M for validation. The model is specifically tuned for the Kandinsky v2.2 diffusion image decoder. Therefore, we use pre-trained OpenCLIP-ViT-G/14222https://huggingface.co/laion/CLIP-ViT-g-14-laion2B-s12B-b42K as the text and vision encoders, ensuring alignment with Kandinsky v2.2 image embeddings. Training is executed on 2 x A100 GPUs, leveraging a per-GPU batch size of 512 and a peak learning rate of 0.00005, across approximately 100,000 iterations, summing up to 74 GPU hours. During inference, the model employs 50 DDIM steps and 7.5 classifier-free guidance for the Kandinsky v2.2 diffusion image generator. Adhering to baseline methodologies, we perform the P-T2I following the baseline papers’ protocols. For $\lambda$-ECLIPSE, target subject pixel regions in reference images are segmented before embedding extraction via the CLIP(vision) encoder. We drop the Canny edge map during inference unless specified explicitly. Unless specified all results (quantitative and qualitative) are without concept-specific additional fine-tuning.
@@
We primarily utilize Dreambench (encompassing 30 unique concepts with 25 prompts per concept) for qualitative and quantitative evaluations using DINO and CLIP-based metrics ruiz2023dreambooth . Due to their limitations, we extend our evaluations on the ConceptBed patel2023conceptbed benchmark (covering 80 diverse imagenet concepts and a total of 33K composite prompts), where we report performance on concept replication, concept, and composition alignment using the Concept Confidence Deviation ($CCD$) metric patel2023conceptbed . We extend Dreambench for multi-subject customization and present the Multibench dataset. Multibench contains about 24 unique concepts and 15 diverse prompts that result in 904 two-subject specific prompts and 1476 three-subject specific prompts. We provide further details about the Multibench in supplementary materials.
Table 2: Quantitative comparisons of different methodologies on Dreambench. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents that we re-benchmark the performance from open-source weights.
Method Base Model Params GPU Hours DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
Textual Inversion SDv1.5 768 1 0.569 0.780 0.255
DreamBooth SDv1.5 0.9B 0.2 0.668 0.803 0.305
Custom Diffusion SDv1.5 57M 0.2 0.643 0.790 0.305
BLIP-Diffusion SDv1.5 0.9B 0.1 0.670 0.805 0.302
@@
Re-Imagen Imagen - - 0.600 0.740 0.270
ELITE SDv1.4 77M 336 0.621 0.771 0.293
Subject-Diffusion SDv1.5 252M - 0.711 0.787 0.293
BLIP-Diffusion* SDv1.5 1.5B 2304 0.603 0.793 0.291
IP-Adapter* SDv1.5 22M 672 0.629 0.827 0.264
IP-Adapter* SDXL 22M 672 0.613 0.810 0.292
[…]
Emu2* SDXL 37B - 0.563 0.765 0.273
$\lambda$-ECLIPSE* Kv2.2 34M 74 0.613 0.783 0.307
Table 3: Quantitative comparisons of different methodologies on ConceptBed. We present results on $CCD~{}(\downarrow)$ across three evaluation categories. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. * represents our benchmarking.
@@
Model Replication (normal-$\downarrow$) Alignment (normal-$\downarrow$) Alignment (normal-$\downarrow$)
@@
4.1 Results & Analysis
Figure 5: Qualitative results categorized by generative capabilities.
@@ 4.1 Results & Analysis
The quantitative assessments detailed in Table 2 and Table 3 focus on the single-concept T2I task, while Table 4 shows the results on multi-concept-driven image generation. For Dreambench and Multibench, we generate and evaluate four images per prompt, reporting average performance on three metrics (DINO, CLIP-I, and CLIP-T). In the case of ConceptBed, we process each of the 33K prompts to generate a single concept image. The results, as depicted in these tables, highlight $\lambda$-ECLIPSE’s superior performance in composition alignment while maintaining competitive concept alignment. Analysis on ConceptBed (Table 3) indicates that $\lambda$-ECLIPSE exhibits a notable proficiency in concept replication, albeit with a marginal trade-off in concept alignment for enhanced composition fidelity. Comparatively, all baselines prioritize concept alignment, often at the expense of composition alignment. While $\lambda$-ECLIPSE improves the CLIP-T while preserving the DINO; achieved with significantly fewer resources.
However, a significant gap remains between finetuning-based and fast P-T2I methods. We further perform concept-specific fine-tuning (as described in Section 3.3). As shown in Table 2, $\lambda$-ECLIPSE outperforms the DreamBooth and BLIP-Diffusion in terms of concept alignment (DINO) while maintaining the performance on composition alignment (CLIP-T). Notably, Multibench results (Table 4) indicate that $\lambda$-ECLIPSE significantly outperforms the Kosmos-G (2B params) and Emu2 (37B params) in terms of CLIP-T while maintaining the DINO performance. Therefore, we can conclude that $\lambda$-ECLIPSE is the most resource-efficient compared, especially when compared to the MLLM-based methods.
Table 4: Quantitative comparisons of different methodologies on Multibench. The Bold and underline represent the metric-specific first and second-ranked methods on each metric, respectively.
Two Subjects Three Subjects
Kosmos-G Emu2 $\lambda$-ECLIPSE Emu2 $\lambda$-ECLIPSE
DINO (normal-$\uparrow$) 0.4549 0.4451 0.4478 0.3168 0.3420
CLIP-I (normal-$\uparrow$) 0.7759 0.7397 0.7409 0.6231 0.6463
CLIP-T (normal-$\uparrow$) 0.2493 0.2673 0.3327 0.2819 0.3469
Table 5: Quantitative results of Canny edge controlled P-T2I of different methodologies on Dreambench. The Bold and underline represent the metric-specific first and second-ranked methods, respectively. Red highlighted numbers indicate the relative percentage drop for concept alignment compared to Table 2.
Method DINO (normal-$\uparrow$) CLIP-I (normal-$\uparrow$) CLIP-T (normal-$\uparrow$)
BLIP-Diffusion* 0.423429.7% 0.7119 0.3152
IP-Adapter* 0.428131.9% 0.7315 0.3034
[…]
@@
In Figure 4, we present a range of single subject-specific images generated by various methodologies including BLIP-Diffusion, IP-Adapter, Kosmos-G, Emu2, and $\lambda$-ECLIPSE. $\lambda$-ECLIPSE demonstrates exemplary proficiency in composition while ensuring concept alignment. In contrast, the baselines often overemphasize reference images or exhibit concept dilution, leading to higher concept alignment but compromised composition. Interestingly, we find that Emu2 can capture the single-subjects but it fails to reproduce them with complex text compositions (as shown in Figure 4). Similarly, Figure 5(a) exhibits $\lambda$-ECLIPSE’s multi-concept generation prowess, in comparison to ZipLoRA (fine-tuning-based approach) along with Kosmos-G and Emu2 (Multimodal LLM-based approaches), underscoring its capability to rival compute-intensive methods. We discuss additional examples and limitations in the appendix. That said, even though $\lambda$-ECLIPSE improves the performance over the baselines, this is still not enough and it signifies the challenges associated with fast multi-concept personalization.
(a) Multi-subject qualitative examples.
(b) Qualitative examples for edge-guided P-T2I.
Figure 6: Qualitative comparison between $\lambda$-ECLIPSE and other baselines.
Canny edge controlled image generation.
As shown in Figure 5(b), the baseline (BLIP-Diffusion) adheres strictly to the imposed edge maps, often at the cost of concept retention (rows 1, 3, and 4). This leads to a large number of unwanted artifacts in the generated images. To further ground this behavior, we first generated images using Stable Diffusion v1.5 for Dreambench prompts without customization then we extracted the Canny edge map and used this edge map to control the subject-driven image generations. At last, we report the performance in Table 5. It can be observed that both baselines IP-Adapter and BLIP-Diffusion drop the DINO score by 30%, which follows the qualitative results. While $\lambda$-ECLIPSE do not follow the Canny edge precisely but preserves the concept alignment and improves the performance relatively by 21%.
@@
Alignment (normal-$\downarrow$) Alignment (normal-$\downarrow$)
Projection loss (i.e. $\lambda$=0.0) 0.394 0.008
 w/ contrastive loss ($\lambda$=0.5) 0.435 -0.043
 w/ contrastive loss ($\lambda$=0.2) 0.402 -0.026
@@
We extend our study to evaluate the individual contributions of different components in $\lambda$-ECLIPSE. Initially, the model’s performance with solely the projection loss (referenced in Eq.1) is assessed. Subsequent experiments involve training $\lambda$-ECLIPSE variants with varying hyperparameters for the contrastive loss, specifically $\lambda$ values of 0.2 and 0.5. A comparative analysis of these baselines is conducted against the fully equipped $\lambda$-ECLIPSE model, which incorporates $\mathcal{L}_{prior}$ (Eq.1) with $\lambda=0.2$ and utilizes Canny edge maps during training. Relying solely on projection loss results in high concept alignment but compromises compositions (Table 6). The contrastive loss variant with $\lambda=0.5$ enhances composition alignment at the expense of concept alignment, whereas $\lambda=0.2$ achieves a more balanced performance. Crucially, the integration of Canny edge maps during training optimally balances both alignments and, specifically, improves the concept alignment. The negative values indicate that the $CCD$ oracle model is highly confident in the generated images.
@@
A key attribute of the CLIP latent space is the ability to perform smooth interpolation between two sets of embeddings. We conducted experiments to demonstrate $\lambda$-ECLIPSE’s ability to learn and replicate this smooth latent space transition. We selected two distinct dog breeds ($<$dog1$>$, $<$dog2$>$) and two types of hats ($<$hat1$>$, $<$hat2$>$) as the concepts. $\lambda$-ECLIPSE was then used to estimate image embeddings for all four possible combinations, each corresponding to the input phrase “a $<$dog$>$ wearing a $<$hat$>$.” Fig. 7 showcases a gradual and seamless transition in the synthesized images from the top left to the bottom right. Unlike current diffusion models, which often exhibit sensitivity to input variations requiring numerous iterations of user interactions for desired outcomes, $\lambda$-ECLIPSE inherits CLIP’s smooth latent space. This not only facilitates progressive changes in the conceptual domain but also extends the model’s utility by enabling personalized multi-subject interpolations.
@@ 5 Conclusion
In this paper, we have introduced a novel diffusion-free methodology for personalized text-to-image (P-T2I) applications, utilizing the latent space of the pre-trained CLIP model with high efficiency. Our $\lambda$-ECLIPSE model, trained on an image-text interleaved dataset, achieves the capability to execute single-concept, multi-concept, and edge-guided controlled P-T2I tasks using a singular model framework, while simultaneously minimizing resource utilization. Notably, $\lambda$-ECLIPSE sets a new benchmark in achieving competitive performance in terms of concept and composition alignment. Furthermore, our research illuminates the potential of $\lambda$-ECLIPSE in exploring and leveraging the smooth latent space. This capability unlocks new avenues for interpolating between multiple concepts and their amalgamation, thereby generating entirely novel concepts. Our findings underscore a promising pathway to improve MLLMs to effectively control the pre-trained diffusion image models without necessitating extra supervision.